There are 20 mg of cholesterol in 8 ounces of yogurt. How many mg of cholesterol are there in 5 ounces of yogurt?

Answer by jim_thompson5910(35256)


(amount of cholesterol)/(amount of yogurt) = (amount of cholesterol)/(amount of yogurt)

(20 mg of cholesterol)/(8 oz of yogurt) = (x mg of cholesterol)/(5 oz of yogurt)

20/8 = x/5

20*5 = 8*x .... cross multiply

100 = 8x

8x = 100

8x/8 = 100/8 ... divided both sides by 8

x = 12.5

There are 12.5 mg of cholesterol in 5 ounces of yogurt.